OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to y = 3x+6 and passes through the point (-10, 2.5)

OpenStudy (zale101):

parallel lines have same slope, agree? :)

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

Making an equation that's parallel to the equation given is just a fancy way of saying I want to move up or down just as fast as that guy, where how fast you're moving is dependent on the slope. In our case the slope is 3 So if we put it into the same equation Y-Yo=m(X-Xo) Y-2.5=3(X+10) Y=3x+12.5

OpenStudy (anonymous):

Yes parallel lines have the same slope

OpenStudy (anonymous):

Hope that clears up your confusion, feel free to ask any more questions to clarify anything

OpenStudy (anonymous):

114 bytes are you a teacher

OpenStudy (anonymous):

Nope, just a student like you.

OpenStudy (anonymous):

Oh, ok. I am taking a ap Geometry

OpenStudy (zale101):

I like your explanation @114Bytes very neat, but we should've directed @One_3234 to solve the answer. Not explaining and then given out the answer :$

OpenStudy (anonymous):

Zale101. I wanted the answer. Not someone teaching me

OpenStudy (anonymous):

My apologies, I tend to think of these problems as plug and chug when it comes to computation what's really more important to me is the illustration of the problem

OpenStudy (anonymous):

Bad thing to say One_3234, I'll keep that in mind next time.

OpenStudy (anonymous):

Ok . Are you a computer or something. Byte 114

OpenStudy (zale101):

Welcome to OS, @One_3234 and @114Bytes. We are here to tutor people and not giving direct answers in the terms of the code of conduct http://openstudy.com/code-of-conduct :) And yes @114Bytes i understand ^_^

OpenStudy (anonymous):

No, not a computer. But studying to become a Computer Engineer.

OpenStudy (anonymous):

I needed the answer. So i can make sure my answer write. and it was

OpenStudy (zale101):

@One_3234 we are real tutors, tutoring online all the users on here are humans for sure

OpenStudy (zale101):

but i'm an alien >:)

OpenStudy (anonymous):

I don;t like getting reported. For something you don't understand about.

OpenStudy (zale101):

How about you show your attempts first, to ensure you wanted to make sure about your answer, and then we will correct you :)

OpenStudy (anonymous):

ok. But I am still reporting you.

OpenStudy (zale101):

You think I reported you?

OpenStudy (zale101):

I never did to let you know :)

OpenStudy (anonymous):

I saw you report. I get updates about who reports. Byte 114 reported me

OpenStudy (anonymous):

ok. bye nice knowing you

OpenStudy (zale101):

what updates?? lol xD

OpenStudy (anonymous):

Lol that's funny because I don't even know how to report people on this site. Haven't bothered to explore the ins and outs yet, care more about the tutor section than the report. I could care less if you wanted to check your answer. It's on you if you just needed the answer for homework.

OpenStudy (anonymous):

Btw, only been here 3 days and it was only today I saw the home page.

OpenStudy (tkhunny):

@One_3234 - You don't seem to have shown any of your own work.